OpenStudy (anonymous):

is x^2+4x-32=0 i got -8,4

OpenStudy (anonymous):

To check whether to have done correctly....Just plug Those Valuse in the eq and u have to get 0 x^2+4x-32 put x= -8 , x=4

OpenStudy (anonymous):

X=4 16 +16 -32 =0...yup..... one of ur soln is correwct....nw check for -8

OpenStudy (anonymous):

do i use 16 what do i use to check it

OpenStudy (anonymous):

Its correct. x^2+4x-32=0 can also be written as (x+8)(x-4)=0. So, (x+8)=0 or (x-4)=0. Therefore x=-8 or 4. Hope it helps.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yep its good. :)